OpenStudy (anonymous):

Two ocean liners leave from the same port in Puerto Rico at 10:00 a.m. One travels at a bearing of N 51o W at 14 miles per hour, and the other travels at a bearing of S 56o W at 17 miles per hour. Approximate the distance between them at noon the same day. Round answer to two decimal places.

OpenStudy (anonymous):

is it 51 degree north of west

OpenStudy (anonymous):

?

OpenStudy (anonymous):

A. 20.33 miles B. 43.25 miles C. 28.20 miles D. 22.72 miles E. 37.19 miles

OpenStudy (anonymous):

that doesnt seem reasonable......because it said round your answer to two decimal points. oh answers!

OpenStudy (anonymous):

Find x and Y component of speed of each ship

OpenStudy (anonymous):

ok so ship a would be 28 miles and shib b would be 34 miles i think

OpenStudy (anonymous):

Here How I would solve it .. Ship 1 travel at 51 degree North of West Using standard coordinate system X component of speed is -cos(51)*14 Y component of speed is sin(51)*14

OpenStudy (anonymous):

This is just ship 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

try this with ship 2

OpenStudy (anonymous):

so you will get speed of two ship

OpenStudy (anonymous):

Multiply x and Y component of each ship seperately to find position then just use distance formula

OpenStudy (anonymous):

ok so ship a is -97.49 ship b is 128.6 ?

OpenStudy (anonymous):

X position of Ship 1 at noon(12-10am) -cos(51)*14 mile per hours * 2hour -17.62 Y position of Ship 2 at noon sin(51)*14*2=21.76

OpenStudy (anonymous):

X position of Ship 2 -cos(56)*17*2=something Y position of Ship 2 -sin(56)*17*2=something Distance Formula \[\sqrt(ship 1 (x) position - ship 2( x) position)^2+ship 1 (y) position - ship 2( y) position)^2)\]

OpenStudy (anonymous):

X position of Ship 2 -cos(56)*17*2=-23.89 Y position of Ship 2 -sin(56)*17*2=17.73

OpenStudy (anonymous):

so then I just plug those into the formula?

OpenStudy (anonymous):

I got 19.01 for x position of ship 2

OpenStudy (anonymous):

hmmm ok you're probably right

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

can you see the distance formula in full

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is basically square root of (xposition of ship 1- xpos ship 2)^2 + (ypos ship 1-y pos ship 2)^2)

OpenStudy (anonymous):

Is it physics or math?

OpenStudy (anonymous):

math trig/pre calc

OpenStudy (anonymous):

Were you able to get answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok I have to go back to the begining and replug in stuff cause i messed something up hold on

OpenStudy (anonymous):

I will do the same thing

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I got 45.73.....?

OpenStudy (anonymous):

The answer is 37.19

OpenStudy (anonymous):

I read that problem wrong it is 51 degree west of north ---- not north of west so I have to switch sin to cos vice versa

OpenStudy (anonymous):

okay I have no clue what i did then

OpenStudy (anonymous):

Ship 1 position X: - sin(51)*14*2 -- A Y: cos(51)*14*2 --B Ship 2 position X: -sin(56)*17*2 -- C Y: -cos(56)*17*2 --D

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sqrt((A-C)^2+(B-D)^2)

OpenStudy (anonymous):

=37.19

OpenStudy (anonymous):

ok thank you!